# MixLM: High-Throughput and Effective LLM Ranking via Text-Embedding Mix-Interaction

### v1

@@ 6.3.5. Inference Optimizations
To evaluate the effectiveness of our inference optimizations for the MixLM framework, we conducted an ablation study measuring maximum throughput (QPS, items per second per GPU), within the same p99 latency budget of ¡=500 ms. The experiments were performed on an NVIDIA H100 80GB HBM3 GPU with CUDA 12.8. The workload consists of a shared prefix of 60 text tokens (search query and user history) and a batch size of 50 candidate items. In different experiment settings, candidate items have different corresponding representations: full item description texts (“raw-text”, 766 tokens/item, 38300 tokens/batch), summarized item description texts (“summarized-text”, 145 tokens/job, 7250 tokens/batch), and (“Mix-LM”, 1 embedding + 1 special token = 2 tokens/job, 100 tokens/batch). Each representation is evaluated without optimizations, with In-Batch Prefix Caching and with Multi-Job Scoring.
Table 11 presents the maximum items throughput per second per GPU, for p99 latency under 500ms. For each configuration, it demonstrates the impact of input compression and inference optimizations. MixLM with In-batch prefix caching achieves the highest throughput (22,000 QPS), contributing to the increase in 10.0x throughput. These results highlight the efficiency gains from compressing job descriptions and optimizing query processing, validated on an H100 GPU.
Table 11. Ablation Study on Inference Optimizations (<500 ms latency)
@@
Summarized, Pruned None 1 650
Multi-Job Scoring 2 100
In-Batch Prefix Cache 2 200
MixLM None 3 000
Multi-Job Scoring 20 000
In-Batch Prefix Cache 22 000
@@ 7. Related Works
7.1. LLMs for Recommender Systems
Recent years have seen a surge of interest in utilizing LLMs for search and recommender system applications. For example, several papers have reported that LLMs’ understanding of natural language makes them powerful relevance judges. Moreover, LLMs are strong ranker which can extract users’ preferences based on their past interactions (relevance4). This has prompted a long line of research, investigating how LLMs can be integrated to existing recommender system, or even used as standalone recommenders (llmrs). For example, (related1) show that LLMs can be successful in solving the cold-start problem in recommender systems, where not much historical data is available for the user. The cold-start problem was further studied by (related5) through the lens of instruction tuning. A similar approach was also used by (related4)(related3) discuss using LLMs to predict user actions and preferences when the model is presented with the past data as a text prompt.
@@ 7.2. LLMs for Semantic Search
More related to our work, there is a long line of research on how LLMs can benefit semantic search applications. A common approach to enhance relevance ranking using LLMs is the bi-encoder approach (relevance-biencoder; relevance-biencoder2), where the query and the item are passed through two separate encoders (or towers) to obtain embedding representations, and the resulting representations are compared to judge the relevance. This approach allows for a computationally efficient implementation, and has been extensively utilized in practice (facebook-ebr). However, this representation-based approach might not be able to capture nuanced semantic clues, due to limited interaction between queries and documents. In response, other work has studied late interaction representation-based models (relevance-colbert).
An alternative approach to using LLMs for semantic search is the cross-encoder approach, where a single model scores the relevance of a query and an item (or multiple items under a list-wise ranking scheme) in a single prompt. Cross-encoder systems with LLM backbones have proved to be strong scorers for ranking systems (sachan2022improving; zhuang2024setwise; zhang2024two; rankvicuna; relevance3). Unfortunately, serving LLM-based cross-encoders as online relevance scorers at an industrial scale can be prohibitively expensive in practice, as scoring numerous pairs of query-item using an LLM would require access to significant computational resources (related-pinterest; relevance-distill; relevance-distill2; rel-distill-reason).
To overcome the computational requirements of LLM-based cross-encoders, other papers have proposed alternative approaches to use LLMs in relevance scoring applications. For example, LLM-as-judge systems make use of LLMs as policy teachers by assigning graded relevance labels and rationales to query–item pairs. This allows for creating training datasets that can be used to distill the knowledge from LLMs to more efficient architectures (related-pinterest; relevance-tencent; relevance-distill3; relevance-distill4; rd-bench; rel-distill-reason). Another approach is LLM-as-augmenter, where LLMs generate paraphrased queries, summarize long documents, and synthesize hard negatives to enrich training data (behdin2025scaling; relevance-neg, and references therein).
7.3. Industrial Application of LLMs for RecSys
In addition to understanding the methodological aspects of MixLM, we also focus on a real-world semantic search that is deployed at an industrial scale. Related to this, (related-pinterest) discuss how an LLM can be used as a relevance judge. However, the LLM-based system is eventually distilled into a non-LLM architecture for online serving. Similar challenges have been reported by (related-ebay) and (related-walmart), where either the model or some of the text features need to be altered for online serving. This further motivates our work to study methods that improve the efficiency of LLM-based recommender systems. In a recent work, (relevance-tencent) study the distillation an LLM-based system into a BERT-type (bert) model. Notably, for most these real-world applications discussed here, to meet efficiency goal, either a non-LLM model has to be deployed, or some features have to be removed. In contrast, in our real-world application, we deploy the LLM online, and all features are consumed by our encoder model to create item embedding representations.
7.4. Mixed Input LLMs for RecSys
Mixed-input LLMs for recommender systems has also been studied by some prior work. (related2) propose to substitute parts of the input text prompt with embeddings available from traditional (non-LLM) embedding systems. In another work, (chen2024hllm) propose to use embeddings to present user’s past interactions, instead of a text prompt. We note that however, our work differs from the aforementioned papers in a few key areas. First, we study a different prediction problem compared to (chen2024hllm; related2) for a relevance scoring system, where nuanced semantic understanding is of high importance. In addition, we present concrete lessons for training our MixLM system, including custom loss and regularization functions that differentiate our work. For example, we show that how an existing pure-text system can be distilled into the mixed-input framework using the custom loss functions we introduce. Finally, (related2)’s analysis is limited to offline benchmarks, while we deploy our mixed input framework online for an A/B test. Moreover, even though chen2024hllm discuss online A/B testing, their hierarchical framework is used offline to generate user and item features, which are then fed to an online recommender system, and therefore, the LLM is not deployed for online inference. This is different from our work where the LLM is used for online ranking. Hence, we present lessons from optimizing the serving infrastructure for such a mixed input system, as well as providing online quality metrics.
Concurrent to our work, (lin2025refrag) studies a mixed input system for Retrieval-Augmented Generation (RAG) applications, which is a different problem from ours. Additionally, we study a real-world deployment of our system, while the analysis of (lin2025refrag) is mostly contained to an offline study. In particular, we present several lessons and insights from serving infrastructure optimizations.
@@ 7.5. Multimodal Large Language Model Training
A common way to build multimodal large language models (MLLMs) is to couple a pretrained text-only LLM with modality-specific encoders via lightweight connectors. Early systems such as Flamingo (alayrac2022flamingo) and Kosmos-1 (huang2023language) showed that treating visual features as additional “tokens” for a transformer backbone enables strong performance on captioning, VQA, and open-ended dialogue, typically by inserting cross-attention layers or training a unified multimodal transformer on large image–text corpora.
Subsequent work has emphasized reusing vision encoders and LLMs, sometime adding a small bridge between them. BLIP-2 (li2023blip) uses a Querying Transformer to align vision features with the LLM embedding space, and LLaVA-style models (liu2023visual) learn a simple projection that maps CLIP features directly into the token embedding space of an LLM, followed by visual instruction tuning. In all these approaches, the key idea is to bypass explicit tokenization of non-text inputs and inject continuous modality-specific embeddings into the LLM, preserving its architecture and linguistic knowledge while adding multimodal capability. Our MixLM follows this paradigm, directly feeding learned embeddings into the language model to replace the tokenizer and embedding layer.
@@ 8. Conclusion
Empirical results from deployment in a large-scale professional social network demonstrate that MixLM delivers substantial system gains—10.0x times improvement in throughput, under same latency budget. MixLM enabled LLM search system’s deployment on all LinkedIn users’ traffic, thereby producing a measurable 0.47% increase in daily active users (DAU). These findings highlight the potential of mixed-input architectures to reshape the design space of LLM-based ranking systems, offering a practical path toward scalable, low-latency, and semantically powerful search in real-world environments.

### v2

@@ 6.3.4. Inference Optimizations
Table 10. Ablation on Inference Optimizations (<<500 ms latency)
@@
Summarized, Pruned None 1,650
Multi-Item Scoring 2,100
In-Batch Prefix Cache 2,200
MixLM None 3,000
Multi-Item Scoring 20,000
In-Batch Prefix Cache 22,000
MixLM combined with in-batch prefix caching yields the highest throughput, confirming the importance of both input compression and shared-prefix amortization for high-QPS LLM ranking.
@@ 7.1. LLMs for Semantic Search
Extensive prior work studies the use of LLMs for semantic search. A widely adopted and computationally efficient approach is the bi-encoder paradigm (Ma et al., 2024; Karpukhin et al., 2020), which independently embeds queries and documents into a shared vector space and is commonly deployed in production (Huang et al., 2020). However, limited query–document interaction can miss fine-grained semantics, motivating late-interaction models that introduce token-level matching (Khattab and Zaharia, 2020).
Cross-encoder models jointly encode queries and documents (or document lists), enabling richer interaction modeling. LLM-based cross-encoders achieve strong ranking performance (Sachan et al., 2022; Zhuang et al., 2024; Zhang et al., 2024; Pradeep et al., 2023; Meng et al., 2025), but their computational cost makes large-scale online serving challenging (Wang et al., 2024; Gao et al., 2020; Hofstätter et al., 2020; Xia et al., 2025).
To address these limitations, prior work leverages LLMs indirectly. The LLM-as-judge paradigm uses LLMs to generate graded relevance labels and rationales to supervise efficient student models (Wang et al., 2024; Ye et al., 2025; Tang et al., 2019; Reddi et al., 2021; Qin et al., 2023; Xia et al., 2025), while LLM-as-augmenter approaches enrich training data via paraphrasing, summarization, or hard-negative generation (Behdin et al., 2025; Li et al., 2024). Our work builds on these directions while enabling direct deployment of LLM-based ranking through model–infrastructure co-design.
7.2. Mixed-Input LLMs for Recommendation Systems
Mixed-input LLMs that combine text tokens with embedding inputs have been explored in recommendation settings. Prior work replaces parts of text prompts with embeddings from traditional models (Zhang et al., 2025b) or represents user interaction histories using embeddings rather than text (Chen et al., 2024). These methods primarily target recommendation or offline feature generation.
In contrast, we focus on relevance ranking, where nuanced semantic understanding is critical. We present a complete training recipe for mixed-input ranking, including custom losses and distillation from pure-text LLMs, and deploy the model directly for online inference with reported A/B test results. We further describe serving-system optimizations required to scale mixed-input ranking models in production.
Concurrent work by Lin et al. (2025) studies mixed-input LLMs for retrieval-augmented generation, a setting distinct from ranking and evaluated primarily offline.
@@ 7.3. Multimodal Large Language Model Training
Multimodal LLMs commonly extend pretrained text models by injecting modality-specific embeddings via lightweight adapters. Early systems such as Flamingo (Alayrac et al., 2022) and Kosmos-1 (Huang et al., 2023) treat non-text inputs as additional tokens to enable strong vision–language performance.
Later work emphasizes minimal architectural changes by aligning pretrained encoders with LLM embedding spaces. BLIP-2 (Li et al., 2023) uses a querying transformer to align vision features, while LLaVA-style models (Liu et al., 2023) project CLIP features directly into the LLM token space. MixLM follows this paradigm by injecting learned embedding tokens to replace long text inputs, adapting multimodal techniques to high-throughput semantic ranking.
@@ 8. Conclusion
Empirical results from deployment in a large-scale professional social network demonstrate that MixLM delivers substantial system gains—10.0x times thorughput compared with summarized-text LLM, 75.9x times thorughput compared with full-text LLM, under same latency budget. MixLM enabled LLM search system’s deployment on all LinkedIn users’ traffic, thereby producing a measurable 0.47% increase in daily active users (DAU). These findings highlight the potential of mixed-input architectures to reshape the design space of LLM-based ranking systems, offering a practical path toward scalable, low-latency, and semantically powerful search in real-world environments.